OpenStudy (anonymous):

y varies directly as the square of x. When x = 3, y = 18. What is the value of y when x = 4?

OpenStudy (cwrw238):

y = kx^2 where k = constant x = 3, y = 18 18 = k* 3^2 so k = 18/9 = 2 so the relation is y = 2x^2 so plug in x=4 into this formula and you'll get value of y

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

yw